# QT-DoG: Quantization-aware Training for Domain Generalization

### v1

@@ 4.3.5 Generality with Vision Transformer
In Table 4, we present the results of quantizing a vision transformer (ERM-ViT, DeiT-small) (Sultana et al., 2022) for domain generalization. We compare the performance of the baseline ERM-ViT to its quantized counterpart on the PACS and Terra Incognita datasets, demonstrating QT-DoG’s effectiveness across different architectures. The results clearly show that QT-DoG also improves the performance of vision transformers. Additionally, we provide results for ResNeXt-50 32x4d in the appendix, following a similar evaluation as in Arpit et al. (2022).
Algorithm Backbone PACS TerraInc Compression
ERM_ViT Deit-Small 84.3 $\pm$ 0.2 43.2 $\pm$ 0.2 -
ERM-SD_ViT Deit-Small 86.3 $\pm$ 0.2 44.3 $\pm$ 0.2 -
ERM_ViT + QT-DoG Deit-Small 86.2 $\pm$ 0.3 45.6 $\pm$ 0.4 4.6x
4.3.6 Visualizations
GradCAM Results. In Figure 5, we present some of the examples [^2] More examples are provided in the appendix. from the PACS dataset and show GradCAM (Gildenblat & contributors, 2021) results in the target domain. We perform four different experiments by considering a different target domain for each run, while utilizing the other domains for training. We use the output from the last convolutional layer of the models with and without quantization. Both models are trained under the same settings as in Gulrajani & Lopez-Paz (2021). For our method, we quantize the model after 2000 iteration and employ 7-bit precision as it provides the best out-of-domain performance.
@@ 4.3.6 Visualizations
Figure 5: GradCAM visualization for ERM (Gulrajani & Lopez-Paz, 2021) and QT-DoG. We show results on the PACS dataset (Li et al., 2017) and consider a different domain as test domain in each run, indicated by the different columns in the figure.
@@
These visualizations evidence that quantization focuses on better regions than ERM, and with a much larger receptive field. In certain cases, ERM does not even focus on the correct image region. It is quite evident that quantization pushes the model to learn more generalized patterns, leading to a model that is less sensitive to the specific details of the training set. These qualitative results confirm the quantitative evidence provided in Table 1.
5 Discussion and Limitations
@@
6 Conclusion
We introduced QT-DoG, a novel generalization strategy based on neural network quantization. Our approach leverages the insight that QAT can find flatter minima in the loss landscape, serving as an effective regularization method to reduce overfitting and enhance the generalization capabilities. We demonstrated both theoretically and empirically that quantization not only improves generalization but also stabilizes the training process. Our extensive experiments across diverse datasets show that incorporating quantization with an optimal bit-width significantly enhances domain generalization, yielding performance comparable to existing methods while reducing the model size. Additionally, we proposed EoQ, a powerful ensembling strategy that addresses the challenges of memory footprint and computational load by creating ensembles of quantized models. EoQ outperforms state-of-the-art methods while being approximately four times smaller than its full-precision ensembling counterparts.
Appendix A Per-Domain Performance Improvement
We also report per-domain performance improvement for PACS (Li et al., 2017) and Terra Incognito (Beery et al., 2018) dataset. We choose the best model based on the validation set and report the results in 5 and 6. The results with quantization correspond to 7 bit-precision and we perform quantization after 2000 steps. Table  5 and 6 show that EoQ is consistently better than the current state-of-the-art methods across domains for different datasets.
@@ Appendix A Per-Domain Performance Improvement
Appendix B Bit Precision Analysis Extended
@@
However, as shown in Table 8, decreasing bit-precision through quantization does not always improve performance above the baseline; after a point, there is a tradeoff between compression and generalization. Specifically, our experiments with 4-bit precision and lower did not yield satisfactory results - see the Table below. Finding the sweet spot for balancing speed and performance can be an interesting research direction. Our results evidence that there exist configurations that can improve both speed and performance.
@@
Appendix C Experiments with larger pre-training datasets
@@
Appendix D In-domain Performance Improvement using Quantization
We further study the in-domain test accuracy of our quantization approach without ensembling on PACS and TerraIncognita datasets.
As (Cha et al., 2021), we split the in-domain datasets into training (60%), validation (20%), and test (20%) sets. We choose the best model based on the validation set and report the results on the test set in Table 10. The results with quantization correspond to 7 bit-precision.
QT-DoG also enhances the in-domain performance. The regularization effect introduced by quantization prevents the model from overfitting to edge cases and pushes it to learn more meaningful and generalizable features, which we also demonstrate in Section 4.3.6. As the training data consists of various domains and the quantization limits the range of weight values, it discourages the model from becoming overly complex and overfitting to the noise in the training data. Therefore, the model is more robust to minor input fluctuations.
@@ Appendix D In-domain Performance Improvement using Quantization
Appendix E Visualization
E.1 More GradCAM Results
In Figure 78910, we present some of the examples from the Terra dataset and show GradCAM (Gildenblat & contributors, 2021) results on the target domain. We use the output from the last convolutional layer of the models with and without quantization for GradCAM. Similar to our experiments on PACS dataset, we perform four different experiments by considering a different target domain for each run, while utilizing the other domains for training. Both models are trained with the similar settings as (Gulrajani & Lopez-Paz, 2021). For quantization method, we quantized the model after 2000 iteration and employ 7 bit-precision as it provides the best out-of-domain performance. Moreover, we present some more examples for PACS dataset in Figure 6
@@ E.1 More GradCAM Results
Appendix F Reproducibility
@@
Every experiment in our work was executed on a single NVIDIA A100, Python 3.8.16, PyTorch 1.10.0, Torchvision 0.11.0, and CUDA 11.1.

### v2

@@ 4.1.3 Different Quantization Methods
4.1.4 Results on Wilds Dataset
We performed experiments with 7 bit quantization on two datasets from the WILDS benchmark (Koh et al., 2021). We utilized the same experimental settings as outlined in the WILDS benchmark repository and incorporated quantization into the training process. The results presented in Table 4.1.1 confirm our findings on Domainbed benchmark. We used the same BERT model (Sanh et al., 2019) as in WILDS (Koh et al., 2021). These results highlight that QT-DoG generalizes well across both architectural variations and input modalities, including text.
Algorithm PACS TerraInc Compression
ERM_ViT 84.3 $\pm$ 0.2 43.2 $\pm$ 0.2 -
ERM-SD_ViT 86.3 $\pm$ 0.2 44.3 $\pm$ 0.2 -
ERM_ViT + QT-DoG 86.2 $\pm$ 0.3 45.6 $\pm$ 0.4 4.6x
@@ 4.1.5 Generality with Vision Transformer
In Table 4.1.4, we present the results of quantizing a vision transformer (ERM-ViT, DeiT-small) (Sultana et al., 2022) for domain generalization. We compare the performance of the baseline ERM-ViT to its quantized counterpart on the PACS and Terra Incognita datasets, demonstrating QT-DoG’s effectiveness across different architectures. The results clearly show that QT-DoG also improves the performance of vision transformers.
Table 6: Performance comparison with CLIP-based methods. We report accuracy on DomainNet, TerraIncognita, and Office datasets, as well as the average performance (AVG). QT-DoG achieves competitive accuracy while offering substantial compression.
Algorithm Backbone DomainNet TerraInc Office Avg. C
ERM CLIP 59.9 $\pm$ 0.1 60.9 $\pm$ 0.2 83.0 $\pm$ 0.1 67.9 -
CLIPood CLIP 63.5 $\pm$ 0.1 60.5 $\pm$ 0.4 87.0 $\pm$ 0.1 70.3 -
QT-DoG CLIP 63.1 $\pm$ 0.2 61.9 $\pm$ 0.3 86.7 $\pm$ 0.2 70.6 4.6x
4.1.6 Bit Precision Analysis
Here, we empirically analyze the effect of different bit-precisions for quantization on the generalization of the model. We perform experiments with four different bit levels and present an analysis in Figure 5 on the PACS (Li et al., 2017) and TerraIncognita (Beery et al., 2018) datasets. We report the test domain accuracy averaged across all domains. For both datasets, 7-bit precision was found to be the optimal bit precision to have the best out-of-domain generalization while maintaining in-domain accuracy. Nonetheless, 8 bits and 6 bits also show improvements, albeit smaller than with 7-bit quantization. These results evidence that, even with a 6 times smaller model, quantization still yields better out-of-domain performance without sacrificing the in-domain accuracy.
@@ 4.1.6 Bit Precision Analysis
Figure 6: GradCAM visualization for ERM (Gulrajani & Lopez-Paz, 2021) and QT-DoG. We show results on the PACS dataset (Li et al., 2017) and consider a different domain as test domain in each run, indicated by the different columns.
4.1.7 Generality with CLIP-Based Methods
We compare QT-DoG with existing CLIP-based domain generalization methods using the ViT-B/16 backbone to maintain consistency in architecture. As shown in the table below, QT-DoG achieves higher average accuracy than both ERM (Gulrajani & Lopez-Paz, 2021) and CLIPood (Shu et al., 2023) across the DomainNet, TerraIncognita, and Office datasets. It provides a 0.3 percent improvement over the best baseline, CLIPood. Importantly, this performance gain comes with a 4.6 times reduction in model size, showing that QT-DoG improves both generalization and efficiency.
4.2 Visualizations
@@
In Figure 6, we present some of the examples [^1] More examples are provided in the appendix. from the PACS dataset and show GradCAM (Gildenblat & contributors, 2021) results in the target domain. We perform four different experiments by considering a different target domain for each run, while utilizing the other domains for training. We use the output from the last convolutional layer of the models with and without quantization. These visualizations evidence that quantization focuses on better regions than ERM, and with a much larger receptive field. In certain cases, ERM does not even focus on the correct image region. It is quite evident that quantization pushes the model to learn more generalized patterns, leading to a model that is less sensitive to the specific details of the training set.
5 Conclusion
We introduced QT-DoG, a novel generalization strategy based on neural network quantization. Our approach leverages the insight that QAT can find flatter minima in the loss landscape, serving as an effective regularization method to reduce overfitting and enhance the generalization capabilities. We empirically demonstrated, supported by analytical insights, that quantization not only enhances generalization but also helps stabilize the training process. Our extensive experiments across diverse datasets show that incorporating quantization with an optimal bit-width significantly enhances domain generalization, yielding performance comparable to existing methods while reducing the model size. Additionally, we proposed EoQ, a powerful ensembling strategy that addresses the challenges of memory footprint and computational load by creating ensembles of quantized models. EoQ outperforms state-of-the-art methods while being approximately four times smaller than its full-precision ensembling counterparts.
Acknowledgements
We thank Soumava Kumar Roy, Chen Zhao, Ahmad Jarrar Khan and Muhammad Zakwan for their help. This project has received funding from the European Union’s Horizon 2020 research and innovation programme under the Marie Skłodowska-Curie grant agreement No. 945363. Moreover, this work was funded in part by the Swiss National Science Foundation and the Swiss Innovation Agency (Innosuisse) via the BRIDGE Discovery grant No. 194729.
Impact Statement
In this paper, we presented our research work aiming to highlight the relation of model quantization for the domain generalization problem setup. In considering the societal impact of our approach, a few concerns emerge, such as performance degradation in extreme-level quantization. Although the proposed approach provides competitive performance compared to previous state-of-the-art models on domain generalization benchmarks, we note that performance slightly varies in extreme quantization cases. Our future work is aimed at a thorough analysis of this aspect in safety-critical applications, such as medical imaging, autonomous driving, or surveillance.
Appendix A Datasets and Metrics
PACS (Li et al., 2017) is a 7 object classification challenge encompassing four domains, with a total of 9,991 samples. It serves to validate our method in smaller-scale settings. VLCS (Fang et al., 2013) poses a 5 object classification problem across four domains. With 10,729 samples, VLCS provides a good benchmark for close Out-of-Distribution (OOD), featuring subtle distribution shifts simulating real-life scenarios. OfficeHome (Venkateswara et al., 2017) comprises a total of 15,588 samples. It presents a 65-way classification challenge featuring everyday objects across four domains. TerraIncognita (Beery et al., 2018) addresses a 10 object classification challenge of animals captured in wildlife cameras, with four domains representing different locations. The dataset contains 24,788 samples, illustrating a realistic use-case where generalization is crucial. DomainNet (Peng et al., 2019) provides a 345 object classification problem spanning six domains. With 586,575 samples, it is one of the largest datasets. Furthermore, we present results on WILDS benchmark datasets to demonstrate the effectiveness of our approach in real-world applications.
Appendix B Implementation Details
We use LSQ (Esser et al., 2020) and INQ (Zhou et al., 2017) for model quantization, with the same configuration as existing quantization methods (Esser et al., 2020; Bhalgat et al., 2020; Dong et al., 2019; Yao et al., 2020; Zhou et al., 2017), where all layers are quantized to lower bit precision except the last one. We quantize the models at 8,000 steps for DomainNet and 2,000 steps for the other datasets. Moreover, each channel in a layer has a different scaling factor $s$.
Appendix C Discussion and Limitations
@@
Appendix D Per-Domain Performance Improvement
We also report per-domain performance improvement for PACS (Li et al., 2017) and Terra Incognito (Beery et al., 2018) dataset. We choose the best model based on the validation set and report the results in 7 and 8. The results with quantization correspond to 7 bit-precision and we perform quantization after 2000 steps. Table  7 and 8 show that EoQ is consistently better than the current state-of-the-art methods across domains for different datasets.
@@ Appendix D Per-Domain Performance Improvement
Appendix E Bit Precision Analysis Extended
@@
However, as shown in Table 10, decreasing bit-precision through quantization does not always improve performance above the baseline; after a point, there is a tradeoff between compression and generalization. Specifically, our experiments with 4-bit precision and lower did not yield satisfactory results. Finding the sweet spot for balancing speed and performance can be an interesting research direction. Our results evidence that there exist configurations that can improve both speed and performance.
@@
Appendix F Experiments with larger pre-training datasets
@@
Appendix G In-domain Performance Improvement using Quantization
We further study the in-domain test accuracy of our quantization approach without ensembling on PACS and TerraIncognita datasets. As (Cha et al., 2021), we split the in-domain datasets into training (60%), validation (20%), and test (20%) sets. We choose the best model based on the validation set and report the results on the test set in Table 12. The results with quantization correspond to 7 bit-precision.
QT-DoG also enhances the in-domain performance. The regularization effect introduced by quantization prevents the model from overfitting to edge cases and pushes it to learn more meaningful and generalizable features, which we also demonstrate in Section 4.2. As the training data consists of various domains and the quantization limits the range of weight values, it discourages the model from becoming overly complex and overfitting to the noise in the training data. Therefore, the model is more robust to minor input fluctuations.
@@ Appendix G In-domain Performance Improvement using Quantization
Appendix H Ablation on Layerwise and Channelwise scale
We conducted an ablation study where we set $s$ at the layer level, rather than on a per-channel basis. We see that Channelwise $s$ can lead to 1.5% accuracy as compared to layerwise $s$. The results of this experiment on the PACS dataset with 7 bit quantization are shown below:
Scale OOD Accuracy
No quantization 84.7 ± 0.5
Layerwise 86.3 ± 0.4
Channelwise 87.8 ± 0.3
Table 13: OOD Accuracy with channelwise vs layerwise Scaling factor for quantization.
Appendix I Ablation on Quantization Steps
We conducted an ablation study on the PACS dataset to identify the optimal number of steps after which quantization should be applied. We perform 7-bit quantization and the results are summarized below:
Quantization Step OOD Accuracy
No quantization 84.7 ± 0.5
1000 86.2 ± 0.4
2000 87.8 ± 0.3
3000 86.9 ± 0.4
4000 85.1 ± 0.3
Table 14: OOD Accuracy across different quantization steps.
Appendix J Adding Uniform Noise to weights
Quantization noise and uniform weight noise share similarities in that both introduce perturbations to the model’s parameters. However, quantization noise specifically arises from the discretization of the weights, which can lead to a more structured form of regularization due to the rounding or truncation during the quantization process. In contrast, uniform weight noise typically adds random perturbations with a uniform distribution, which may not exhibit the same structured regularization properties.
In Table 15, we provide the results of our ablation study on the PACS dataset with uniform noise with different minimum and maximum value:
Noise OOD Accuracy
No noise 84.7 ± 0.5
Uniform(-0.0001, 0.0001) 82.9 ± 0.6
Uniform(-0.00005, 0.00005) 83.8 ± 0.5
Uniform(-0.00001, 0.00001) 85.1 ± 0.4
Uniform(-0.000005, 0.000005) 85.6 ± 0.3
Table 15: OOD Accuracy under different noise levels
Appendix K Visualization
More GradCAM Results
In Figure 9101112, we present some of the examples from the Terra dataset and show GradCAM (Gildenblat & contributors, 2021) results on the target domain. We use the output from the last convolutional layer of the models with and without quantization for GradCAM. Similar to our experiments on PACS dataset, we perform four different experiments by considering a different target domain for each run, while utilizing the other domains for training. Both models are trained with the similar settings as (Gulrajani & Lopez-Paz, 2021). For quantization method, we quantized the model after 2000 iteration and employ 7 bit-precision as it provides the best out-of-domain performance. Moreover, we present some more examples for PACS dataset in Figure 7.
@@ Appendix I Ablation on Quantization Steps
Appendix L Quantization Noise follows Uniform Distribution
It is a well established fact in the literature that the distribution of quantization error can be modeled as uniform distribution (Boncelet, 2009; Gray & Neuhoff, 1998). For uniform data distribution, it only needs to hold within bounds of quantization bins as in this case half of the samples will be floored while the other half will be ceiled. As the the distribution of data is uniform with a flat pdf, the error introduced by the rounding operation will also have the same distribution.
In many scenarios, however, the distribution of network weights is not uniform. Therefore, we extend this argument to more general data distributions. If the data follows a symmetric distribution with a reasonably smooth pdf, we can approximate the pdf of this distribution by a piecewise linear function, interpolating between the boundaries of the quantization bins. This approximation becomes more accurate as we increase the bit-width, which in turn increases the number of quantization bins. Although these linear pdfs within the quantization bins are not flat, the symmetry of the distribution ensures that there exists another bin with a similar but negatively sloped pdf. Thus, when considered as a whole, the quantization error becomes uniform, i.e $\epsilon\sim U\left[-\frac{s}{2},\frac{s}{2}\right]$.
In order to model the quantization error as a uniform error, we need to have a symmetric weight distribution. Deep neural networks are overparameterized models that can have multiple optimal solutions. To guide the models toward specific distributions, weight priors are typically applied in the form of regularizers. These priors lead to trained weights that correspond to the specified distribution. L1 regularization encourages sparsity and promotes a Laplace distribution for the weights, which is symmetric and has sharper peaks around zero. On the other hand, L2 regularization promotes a Gaussian distribution, which is also symmetric but follows a smooth bell-shaped curve with heavier tails compared to the Laplace distribution.
These regularization techniques result in weight distributions that are symmetric and smooth, satisfying the necessary condition for quantization error to be modeled as a uniform distribution, where the quantization noise/error exhibits a very low KL divergence with respect to a uniform distribution.
We plot various layers’ weight distributions along with the quantization noise in Figure 8 and demonstrate that weights generally exhibit a symmetric and smooth distribution. Additionally, the quantization noise/error shows a very low Kullback-Leibler (KL) divergence (Kullback & Leibler, 1951) with respect to a uniform distribution having the same minimum and maximum values. To assess symmetry, we employ reverse-order and forward-order medians, ensuring a balanced representation of the weight distribution. The reverse-order median captures the statistical properties of the lower half, while the forward-order median does the same for the upper half, providing a more robust symmetry evaluation.
Figure 8: Distribution of Quantization Noise and Weights. We plot the weights(left), quantization noise(middle) and symmetry(right) of various layers in ResNet-50 model. We found KL-divergence to be [Top {0.0061, 0.004, 00069, 0.00054} Bottom] between quantization noise and uniform distribution with same minimum and maximum value.
Appendix M Reproducibility
@@
Every experiment in our work was executed on a single NVIDIA A100, Python 3.8.16, PyTorch 1.10.0, Torchvision 0.11.0, and CUDA 12.1.